# Smart Glasses for CVI: Co-Designing Extended Reality Solutions to Support Environmental Perception by People with Cerebral Visual Impairment

### v1

@@ 7.2. What are the key considerations and challenges for smart glasses to be an effective assistive tool?
CVI is not a single condition but an umbrella term encompassing a wide range of brain-based visual processing difficulties (Dutton, 2003; Roman-Lantzy, 2007; Lueck et al., 2019). People with CVI may present with varied impairments, including both lower and higher-level visual processing impairments (Philip and Dutton, 2014; Lueck et al., 2019). In the following sections, we illustrate the need for hyper-personalisation informed design by considering user-specific visual conditions and context-aware adaptations.
@@
Even within our small co-designer group, solution preferences varied due to differences in visual impairments. For example, overlays for locating objects had to avoid the right visual field for Nicola due to her hemianopsia. Similarly, determining what constitutes a ‘salient’ feature depends on individual attention profiles—for instance, a tendency to miss blue-coloured objects could warrant targeted visual markers only on them. The onset and developmental history of CVI also shaped interaction preferences. Dijana benefited from the face-to-face conversation aid, while Nicola had already developed robust compensatory strategies, reducing the perceived value of it. These observations underscore the limitations of one-size-fits-all approaches and highlight the need for highly adaptive systems (Aflatoony and Kolarić, 2022; Wey, 2004; Hodson et al., 2024).
@@
Both co-designers expressed interest in integrating physiological indicators—such as heart rate variability (Rowe et al., 1998), eye movements(Jacob, 1990), or pupil dilation (Pomplun and Sunkara, 2019)—to inform real-time adjustments. These bio-signals are increasingly recognised as reliable proxies for cognitive load and emotional state (Schultz and Maedche, 2023; Schmidt et al., 2018), offering a promising pathway towards truly adaptive and responsive assistive experiences. This points to the value of designing adaptive systems that use real-time signals—not just from the environment but also from the user—to modulate assistive strategies dynamically.
@@ Acknowledgements.
We are grateful to the Monash AI Institute for providing the funding that made this project possible.

### v3

@@ 7.2. What are the key considerations and challenges for smart glasses to be an effective assistive tool?
CVI is not a single condition but an umbrella term encompassing a wide range of brain-based visual processing difficulties (Dutton, 2003; Roman‑Lantzy, 2007; Lueck et al., 2019). People with CVI may present with varied impairments, including both lower and higher-level visual processing impairments (Philip and Dutton, 2014; Lueck et al., 2019). In the following sections, we illustrate the need for hyper-personalisation informed design by considering user-specific visual conditions and context-aware adaptations.
@@
Even within our small co-designer group, solution preferences varied due to differences in visual impairments. For example, overlays for locating objects had to avoid the right visual field for Nicola due to her hemianopsia. Similarly, determining what constitutes a ‘salient’ feature depends on individual attention profiles—for instance, a tendency to miss blue-coloured objects could warrant targeted visual markers only on them. The onset and developmental history of CVI also shaped interaction preferences. Dijana benefited from the face-to-face conversation aid, while Nicola had already developed robust compensatory strategies, reducing the perceived value of it. These observations underscore the limitations of one-size-fits-all approaches and highlight the need for highly adaptive systems (Aflatoony and Kolarić, 2022; Wey, 2004; Hodson et al., 2025).
@@
Both co-designers expressed interest in integrating physiological indicators—such as heart rate variability (Rowe et al., 1998), eye movements(Jacob, 1990), or pupil dilation (Pomplun and Sunkara, 2003)—to inform real-time adjustments. These bio-signals are increasingly recognised as reliable proxies for cognitive load and emotional state (Schultz and Maedche, 2023; Schmidt et al., 2018), offering a promising pathway towards truly adaptive and responsive assistive experiences. This points to the value of designing adaptive systems that use real-time signals—not just from the environment but also from the user—to modulate assistive strategies dynamically.
@@ Acknowledgements.
We are grateful to the Monash Data Futures Institute for providing the funding that made this project possible.